# Unique Security and Privacy Threats of Large Language Model: A Comprehensive Survey

### v1

@@ 9.2. Watermarks for LLMs
LLMs have two copyright issues that affect privacy and security. First, LLMs cost massive corpora and computational resources, thus, the model is considered the owner’s property. Second, to clarify the responsibility for generated content, researchers consider copyright issues of the content produced by LLMs, which can mitigate the misuse of LLMs. For model copyright, He et al. (He et al., 2022) proposed a watermark injection scheme that does not modify model parameters. They applied a word replacement strategy to add watermarks to the output. However, Zhao et al. (Zhao et al., 2023a) found that knowledge distillation could remove such watermarks. Therefore, Li et al. (Li et al., 2024c) used backdoor attacks to inject watermarks into customized LLMs. Subsequently, the model owner can efficiently complete verification by checking the backdoor effect. Similarly, Liu et al. (Liu et al., 2023c) embedded the trigger into private texts to trace the data intelligence.
@@
For output copyright, it aims to identify whether a specific LLM generates a given text. Some researchers focus on injecting unique and imperceptible watermarks into generated texts without affecting their semantics. Sato et al. (Sato et al., 2023) modified the output format without changing its content. For instance, they used different types of spaces to replace standard spaces, which could be implemented with minimal change. Yoo et al. (Yoo et al., 2023) found that paraphrasing would remove such watermarks, hence they designed a robust multi-bit watermarking scheme. Specifically, they identified features invariant to minor perturbations and used a BERT-based mask language model to embed watermarks. Zhang et al. (Zhang et al., 2023a) mixed the generated text with binary signatures in the feature space. They used the Gumbel-Softmax function during the encoding phase to transform the generated dense distribution into a sparse distribution. It can significantly enhance the coherence and semantic integrity of watermarked texts.
@@
Other researchers focus on the generation process, directly returning watermarked texts instead of modifying output results. At first, Kirchenbauer et al. (Kirchenbauer et al., 2023) embedded watermarks by manipulating the log of LLMs. They divided the vocabulary into red and green lists based on a random seed, encouraging the LLM to choose tokens from the green list. Then, users who know the partition mode can implement the verification by calculating the number of green tokens in the generated text. However, Hu et al. (Hu et al., 2023) found a trade-off between watermark strength and output perplexity. Thus, they embedded watermarks by manipulating the probability distribution of the next token, preferring specific tokens without changing the text semantics. Similarly, Liu et al. (Liu et al., 2023e) proposed a semantic-invariant watermark scheme, generating watermark logits based on the semantics of all preceding tokens instead of simply token IDs. It maintains the semantics of generated texts, successfully resisting synonym replacement attacks and text paraphrasing attacks. Additionally, they found that these watermark schemes used the same key for generation and verification, which posed a forgery risk in the public verification scenario. Inspired by this observation, literature (Liu et al., 2023d) improved the watermarking scheme proposed by Kirchenbauer et al. (Kirchenbauer et al., 2023). They adopted two neural networks for watermark generation and verification, improving the accuracy of verification and the non-forgery of watermarks. Kuditipudi et al. (Kuditipudi et al., 2023) found that this scheme caused the LLM to generate identical watermarked texts for the same prompt, affecting the LLM’s utility. They used longer pseudorandom number sequences to solve this issue and randomly set the initial position during watermark embedding. Subsequently, Hou et al. (Hou et al., 2023) focused on sentence-level sampling and injected watermarks into it, making the watermark resistant to text paraphrasing attacks.
@@
As users increasingly focus on personal privacy protection, especially under legal frameworks such as the GDPR, the right to forget has become a critical problem. Upon receiving a forgetting request, the model owner can leverage machine unlearning to protect their data from unauthorized use. This technology effectively removes the influence of specific user data. Machine unlearning for LLMs faces three challenges compared to traditional small-scale language models. First, LLMs are usually trained on a large amount of data. However, the contribution of a single sample point is not significant, making it challenging to exact unlearning for specific data. Second, machine unlearning will affect the performance of LLMs, requiring a balance between protecting user privacy and maintaining model performance. Third, concerning LLMs, there needs to be more research on verifying the effectiveness of forgetting. Currently, Yao et al. (Yao et al., 2023) studied machine unlearning for LLMs. They found that the memory capabilities of LLMs far exceeded those of traditional models, running multiple unlearning operations to eliminate the impact of specific data. They also identified the issue of catastrophic forgetting, which affects model performance. Meanwhile, Eldan et al. (Eldan and Russinovich, 2023) addressed copyright issues in corpora, replacing ‘Harry Potter’ with other concepts. They made the target LLM forget content related to ‘Harry Potter’. In terms of verification, Viswanath et al. (Viswanath et al., 2024) evaluated the model’s response to forgot samples. In addition, they explored some verification methods, such as data extraction attacks. Despite the many challenges machine unlearning and verification face, research in this area is crucial for improving the transparency of LLMs. Many countries have introduced regulations on AI models, aiming to improve the transparency of AI-based applications. With transparency, it is easier for researchers to understand the inference of LLMs. Improving transparency will be conducive to further research on LLMs’ privacy and security. Mitchell et al. (Mitchell et al., 2019) improve the transparency of models by establishing model cards. Then, Wahle et al. (Wahle et al., 2023) proved that model cards were suitable for LLMs. It helps users better understand model performance and potential risks to train and deploy LLMs more safely.
@@ 10. Conclusion
This paper provides a comprehensive overview of privacy risks, security risks, and defenses, focusing on LLMs. Specifically, we introduce various privacy and security attacks, including training-time attacks and inference-time attacks, along with their respective subcategories. We also discuss general and specific protection strategies for privacy and security risks. Furthermore, we present LLMs’ evaluation and provide details on the datasets and evaluation metrics regarding utility and robustness. Concerning LLM privacy and security aspects, this paper gives a comprehensive view of attacks and defenses; we believe it is a significant contribution to help researchers understand LLMs’ issues, making them be applied to more fields.

### v2

@@ 9.2. Watermarks for LLMs
LLMs have two copyright issues that affect privacy and security. First, LLMs cost massive corpora and computational resources. Thus, the model is considered the property of the owner. Second, to clarify the responsibility for generated content, researchers consider copyright issues of the content produced by LLMs, which can mitigate the misuse of LLMs. For model copyright, He et al. (He et al., 2022) proposed a watermark injection scheme that does not modify model parameters. They applied a word replacement strategy to add watermarks to the output. However, Zhao et al. (Zhao et al., 2023a) found that knowledge distillation could remove such watermarks. Therefore, Li et al. (Li et al., 2024c) used backdoor attacks to inject watermarks into customized LLMs. Subsequently, the model owner can efficiently complete verification by checking the backdoor effect. Similarly, Liu et al. (Liu et al., 2023c) embedded the trigger into private texts to trace the data intelligence.
@@
For output copyright, it aims to identify whether a given text is generated by a specific LLM. Some researchers focus on injecting unique and imperceptible watermarks into generated texts without affecting their semantics. Sato et al. (Sato et al., 2023) modified the output format without changing its content. For instance, they used different types of spaces to replace standard spaces, which could be implemented with minimal change. Yoo et al. (Yoo et al., 2023) found that paraphrasing would remove such watermarks, hence, they designed a robust multi-bit watermarking scheme. Specifically, they identified features invariant to minor perturbations and used a BERT-based mask language model to embed watermarks. Zhang et al. (Zhang et al., 2023a) mixed the generated text with binary signatures in the feature space. They used the Gumbel-Softmax function during the encoding phase to transform the generated dense distribution into a sparse distribution. It can significantly enhance the coherence and semantic integrity of watermarked texts.
@@
Other researchers focus on the generation process, directly returning watermarked texts instead of modifying output results. At first, Kirchenbauer et al. (Kirchenbauer et al., 2023) embedded watermarks by manipulating the log of LLMs. They divided the vocabulary into red and green lists based on a random seed, encouraging the LLM to choose tokens from the green list. Then, users who know the partition mode can implement the verification by calculating the number of green tokens in the generated text. Then, literature (Liu et al., 2023d) improved the watermark scheme proposed by Kirchenbauer et al. (Kirchenbauer et al., 2023). They adopted two neural networks for watermark generation and verification, improving the accuracy of verification and the non-forgery of watermarks. Kuditipudi et al. (Kuditipudi et al., 2023) found that this scheme caused the LLM to generate identical watermarked texts for the same prompt, affecting the LLM’s utility. They used longer pseudorandom number sequences to solve this issue and randomly set the initial position during watermark embedding.
@@
As users increasingly focus on personal privacy protection, especially under legal frameworks such as the GDPR, the right to forget has become a critical problem. Upon receiving a forgetting request, the model owner can leverage machine unlearning to protect the requested data from unauthorized use. This technology effectively removes the influence of specific user data. Compared to small-scale language models, machine unlearning for LLMs faces three challenges. First, LLMs are usually trained on a large amount of data. However, the contribution of a single sample point is not significant, making it challenging to exact unlearning for specific data. Second, machine unlearning will affect the performance of LLMs, requiring a balance between protecting user privacy and maintaining model performance. Third, concerning LLMs, there needs to be more research on verifying the effectiveness of machine unlearning.
Currently, Yao et al. (Yao et al., 2023) studied machine unlearning for LLMs. They found that the memory capabilities of LLMs far exceeded those of small-scale models, running multiple unlearning operations to eliminate the impact of specific data. They also identified the issue of catastrophic forgetting, which affects model performance. Meanwhile, Eldan et al. (Eldan and Russinovich, 2023) addressed copyright issues in corpora, replacing ‘Harry Potter’ with other concepts. They made the target LLM forget content related to ‘Harry Potter’. In terms of verification, Viswanath et al. (Viswanath et al., 2024) evaluated the model’s response to forgot samples. In addition, they explored some verification schemes, such as data extraction attacks. Despite facing many challenges in machine unlearning and verification, research in this area is crucial for improving the transparency of LLMs.
@@ 10. Conclusion
LLMs have become a strong driving force in revolutionizing a wide range of applications. However, these applications have revealed various vulnerabilities due to the privacy and security threats LLMs face. This has drawn significant academic and industrial attention. Moreover, the threats faced by LLMs differ significantly from those encountered by traditional models. In this paper, we provide a comprehensive and novel overview of privacy and security risks, and discuss their countermeasures, highlighting unique threats for LLMs. By analyzing the life cycle of LLMs, we consider five unique threat scenarios: pre-training LLMs, fine-tuning LLMs, using the RAG system, deploying LLMs, and deploying LLM-based agents. Using these scenarios as a taxonomy, we explore various threat models of LLMs. For each threat model, we highlight unique risks to LLMs from privacy and security perspectives. Additionally, we describe common risks to all models and adapt them to LLMs. Regarding the characteristics of each risk, we provide potential countermeasures. Finally, we integrate LLMs with three additional scenarios related to privacy and security threats, such as federated learning, machine unlearning, and watermarking, meriting future research. In summary, we believe this survey makes a significant contribution to helping researchers understand LLMs’ safety threats, enabling their application to more fields.